OpenStudy (anonymous):

a tennis player makes a successful first serve 48% of the time. if she serves 9 times, what is the probability that she gets exactly 3 first serves?

OpenStudy (anonymous):

This Can be done with the binomial probability distribution here p = 0.48 and q = 1-p = 1-0.48 = 0.52 and n = 9 and we have to find P(x=3)